OpenStudy (anonymous):

write the equation of the line that passes through (-2,1) and is perpendicular to the line y=1/2x+1

OpenStudy (anonymous):

Lines that are perpendicular have opposite slopes from each other. so the slope of the line would be -2/1 or just -2

OpenStudy (anonymous):

Because the slope of the line its perpendicular to is 1/2. Change the sign and flip the fraction and its -2/1 or -2

OpenStudy (anonymous):

see how i got that?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

and how who u write that at the end

OpenStudy (anonymous):

the line you were given is y=1/2x+1. thats in the form y=mx+b where m = slope. you want to find the line that is perpendicular to that. lines that are perpendicular to each other have opposite slope. opposite of 1/2 is -2/1

OpenStudy (anonymous):

Now that we have the slope and the point you were given. we can do point slope form and then simplify to get slope intercept form

OpenStudy (anonymous):

 lets do this